OpenStudy (anonymous):

Express 65 minutes as an improper fraction of an hour in simplest form(one hour is 60 minutes). Is the answer A. seven sixths B. thirteen twelfths C. one twentieth or D. sixty five sixtieths ??? Please help! Reply is greatly appreciated!!

OpenStudy (michele_laino):

hint: we can write this: \[\frac{{65}}{{60}} = \frac{{60 + 5}}{{60}} = \frac{{60}}{{60}} + \frac{5}{{60}} = 1 + \frac{5}{{60}}\] please simplify the fraction \(5/60\) at the right side